# Universal Approximation Theory: The basic theory for large language models

### v1

@@ Abstract
Language models have emerged as a critical area of focus in artificial intelligence, particularly with the introduction of groundbreaking innovations like ChatGPT. Large-scale Transformer networks have quickly become the leading approach for advancing natural language processing algorithms. Built on the Transformer architecture, these models enable interactions that closely mimic human communication and, equipped with extensive knowledge, can even assist in guiding human tasks. Despite their impressive capabilities and growing complexity, a key question remains – the theoretical foundations of large language models (LLMs). What is it about the Transformer architecture that makes it so effective for powering intelligent language applications, such as translation and coding? What underlies LLMs’ ability for In-Context Learning (ICL)? How does the LoRA scheme enhance the fine-tuning of LLMs? And what supports the practicality of pruning LLMs? To address these critical questions and explore the technological strategies within LLMs, we leverage the Universal Approximation Theory (UAT) to offer a theoretical backdrop, shedding light on the mechanisms that underpin these advancements.
Universal Approximation Theory: The basic theory for large language models
Wei Wang1, Qing Li1, 1The Hong Kong Polytechnic University, weiuat.wang@connect.polyu.hk, qing-prof.li@polyu.edu.hk
@@ 1 Introduction
In recent years, the rapid emergence of LLMs in the field of artificial intelligence has undoubtedly become one of the most notable advancements within the domain. The core allure of these models stems from their extraordinary capabilities in language processing. Language, as a unique crystallization of human intelligence, serves not only as the external reflection of thought but also as the bridge for communication, the cornerstone for the dissemination of knowledge, and the continuation of civilization, profoundly shaping the identity of humans as a unique species. Thus, endowing machines with the ability to understand and generate language marks a significant leap towards the realization of true artificial intelligence. The emergence of models such as the ChatGPT Radford and Narasimhan (2018); Brown et al. (2020); Achiam et al. (2023), the Llama Touvron et al. (2023), and the PaLM Chowdhery et al. (2023) vividly demonstrates this point.
@@
A distinctive feature of LLMs is their immense parameter size Achiam et al. (2023); Touvron et al. (2023); Chowdhery et al. (2023); Chen et al. (2021); Zeng et al. (2022), often amounting to hundreds of billions or even trillions (for instance, GPT-3’s Brown et al. (2020) 175 billion parameters and PaLM’s Chowdhery et al. (2023) 540 billion parameters). This vast parameter scale lays the foundation for their exceptional language processing capabilities and enables them to exhibit almost human-like traits, such as ICL  Brown et al. (2020); Dong et al. (2022), instruction following Sanh et al. (2021); Ouyang et al. (2022); Wei et al. (2021), and multi-step reasoning Wei et al. (2022). Notably, these colossal models are predominantly trained by tech giants like Google and Microsoft using large-scale GPU clusters Zhao et al. (2023), sparking a research fervor on how to efficiently fine-tune them with limited GPU resources. The advent of Lora Hu et al. (2021) fine-tuning technology has provided an effective pathway for this, allowing for the fine-tuning of large models under resource constraints without the need to adjust all parameters of the original model comprehensively. Moreover, model pruning techniques Sun et al. (2023); Ma et al. (2023) are crucial for deploying large models in resource-constrained environments, aiming to reduce the model size for operation on smaller devices. Faced with the challenge of processing long texts, such as generating summaries or answering questions based on extensive documents—which traditionally requires substantial computational resources—technologies like LongLora Chen et al. (2023) have been developed to tackle the difficulties of processing long contextual texts, further expanding the application boundaries of LLMs. Today, LLMs possess a diverse range of functionalities, from translation and text summarization to automatic code generation, demonstrating their versatility.
@@
Despite LLMs being on a fast track of development towards higher intelligence and reliability, the theoretical foundation behind them remains largely unexplored, shrouded in clouds of uncertainty. The scientific community is actively seeking to uncover the inner mechanisms behind their powerful capabilities, including analyses of ICL mechanism Xie et al. (2021); Min et al. (2022), among others. Against this backdrop, we establish the UAT Cybenko (2007); Hornik et al. (1989) as the mathematical essence of LLMS, and utilize the UAT lens to clarify pivotal technologies and phenomena within contemporary LLMs, aiming to illuminate this central theoretical conundrum from a fresh perspective. Our contributions are as follows:
@@
  • We offer insights into the future development of LLMs.
@@ 2 The Universal Approximation Theory
Up to now, the UAT (Cybenko, 2007) is the most widely recognized fundamental theory in deep learning. However, this theorem is only applicable to the simplest form of neural networks, the multilayer perceptron (MLP) (Cybenko, 2007; Popescu et al., 2009). Due to the increased complexity of Transformer networks, they cannot be mathematically expressed in the same form as UAT. Thus, the theory has not yet been extended to Transformer networks. Our aim in this paper is to unify Transformer networks under the framework of UAT, thereby standardising their mathematical representation. Before unifying their mathematical forms, we provide a brief overview of the theorem, which was initially proposed by Cybenko (2007). This theorem encompasses numerous conclusions and proof details. Although it has been further developed, its fundamental mathematical form remains unchanged. Therefore, this paper explains the theory based on the UAT form presented by Cybenko (2007). Theorem 2 from Cybenko (2007) states that if $\sigma$ is any continuous sigmoidal function, then finite sums of the following form:
@@
This implies that, when $N$ is sufficiently large, a neural network can approximate any continuous function on a closed interval. Hornik et al. (1989) further demonstrates that multilayer feedforward networks also conform to the UAT, capable of approximating arbitrary Borel measurable functions. Observing Equation (1), where the function $G(\mathbf{x})$ yields a scalar output in $\mathbb{R}$, the scenario expands naturally when $G(\mathbf{x})$ maps to $\mathbb{R}^{m}$, requiring the approximation in each dimension. It becomes evident that to accommodate this multidimensional output, a simple extension to Equation (1) suffices: the transformation matrix $\mathbf{W}_{j}$ is revised to the space $\mathbb{R}^{n\times m}$, the bias term $\theta_{j}$ is recast as a vector in $\mathbb{R}^{m}$, and $\alpha_{j}$ is reshaped into a diagonal vector. Nevertheless, in these demonstrations, the theorem does not straightforwardly apply to Transformer architectures. However, if we succeed in reformulating both the Linear ($\mathbf{x}|\mathbf{W}$) and MHA ($\mathbf{x}|\mathbf{W}$) components into a unified representation, $\mathbf{W}^{\prime}\mathbf{x}^{\prime}$, where $\mathbf{W}$ and $\mathbf{x}$ represent the parameters and inputs for each component respectively, and $\mathbf{W}^{\prime}$ is derived from $\mathbf{W}$ while $\mathbf{x}^{\prime}$ signifies a column vector (for details, refer to Section 3.1). It is easy to prove that multi-layer Transformer networks are concrete embodiments of UAT.
@@ 3.2 Diamond Matrix Multiplication
While our goal is to transform Linear and MHA modules into a matrix-vector format, this process is complex. In order to describe this transformation process clearly, we propose a new computational method called the Diamond Multiplication Method, denoted by the symbol $\diamond$.
@@
$=\left(\begin{array}[]{c}w_{1}x_{1}+w_{31}*x_{2}+\cdots w_{m,1}x_{m}\\ w_{1,2}x_{1}+w_{2,2}x_{2}+\cdots w_{m,2}x_{m}\\ \vdots\\ w_{1,n}x_{1}+w_{2,n}x_{2}+\cdots w_{mn}x_{m}\end{array}\right)$
Let $\mathbf{W}\in\mathbb{R}^{(m,n)}$ and $\mathbf{x}\in\mathbb{R}^{(m,1)}$, the Diamond Matrix Multiplication is defined as $\mathbf{y}=\mathbf{W}\diamond\mathbf{x}$, where $\mathbf{y}\in\mathbb{R}^{(m,1)}$. The computation procedure involves element-wise multiplication and summation of corresponding elements from left to right across the columns of matrix $\mathbf{W}$ with vector $\mathbf{x}$. The sum of the multiplication of the $i$th column of $\mathbf{W}$ with $\mathbf{x}$ is assigned as the $i$th element of $\mathbf{y}$. A detailed calculation process is provided in Eq. ((4)). The computation process for $\mathbf{y}=\mathbf{x}\diamond\mathbf{W}$ is identical to that of $\mathbf{W}\diamond\mathbf{x}$, implying that $\mathbf{W}\diamond\mathbf{x}=\mathbf{x}\diamond\mathbf{W}$. However, it should be noted that the Diamond Matrix Multiplication does not possess the properties of associativity and distributivity.
@@ 3.3 Matrix-Vector Method for Linear
Here, $\mathbf{x}_{i}\in\mathbb{R}^{(N,M)}$ and $\mathbf{x}_{i+1}\in\mathbb{R}^{(N,M)}$ represent the input and output of layer $i$, respectively, while $\mathbf{W}_{i}\in\mathbb{R}^{(N,N)}$ represents the parameters of layer $i$. $\mathbf{x}_{i}^{\prime}$, $\mathbf{x}_{i+1}^{\prime}$, and $\mathbf{W}_{i}^{\prime}$ are generated based on $\mathbf{x}^{i}$, $\mathbf{x}^{i+1}$, and $\mathbf{W}_{i}$ using the Matrix-Vector Method. The derivation of the general form can be found in Appendix B.1.
@@
Figure 4: The process of transforming $\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}}_{8})$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{WH}^{\prime}\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$.
Figure 5: This diagram illustrates the process of transforming $\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}})\mathbf{W}_{O}$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{WHO}^{\prime}\diamond\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, different colors in the diagram correspond to different variables.
@@ 3.4 Matrix-Vector Method for MHA
$=$ $\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}$
@@
Here, $h$ represents the number of attention heads, and the input $\mathbf{x}\in\mathbb{R}^{(N,M)}$ is divided into $\mathbf{x}_{1},...,\mathbf{x}_{h}$ based on $h$. The parameters $\mathbf{W}_{Qi},\mathbf{W}_{Ki}$, and $\mathbf{W}_{Vi}$ correspond to $\mathbf{x}_{i}$. The whole process of MHA can be represented in Figure 3. Figure 3.a represnets that the input $\mathbf{x}$ is split into $\mathbf{x}_{1}\cdots\mathbf{x}_{8}$ based on the number of heads. Figure 3.b represents $\mathbf{x}_{i}\mathbf{W}_{Qi}=\mathbf{Q}_{i},\mathbf{x}_{i}\mathbf{W}_{Ki}=\mathbf{K}_{i},\mathbf{x}_{i}\mathbf{W}_{Vi}=\mathbf{V}_{i}$. Figure 3c, d, e represent the process of $\operatorname{softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{T}}{\sqrt{M}}\right)\mathbf{V}_{i}$ and we omit $\sqrt{M}$. Figure 3.f represents $\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}$. Figure 3.g represents the whole process by a matrix multiplication of $(\mathbf{W}^{\prime}_{HVO})^{T}\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, where $\mathbf{W}^{\prime}_{HVO}$ is generated based $\mathbf{H}_{1}\cdots\mathbf{H}_{8}$, $\mathbf{W}_{V1}\cdots\mathbf{W}_{V8}$ and $\mathbf{W}_{O}$. It means that we could represent the whole complex MHA into a matrix multiplication. Next, we give the proof of this process.
@@
Our objective is to express the MHA as $(\mathbf{W}^{\prime})^{T}_{HVO}\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$. Prior to transforming the MHA into the matrix-vector form, we need to conduct a comprehensive analysis and clearly define the research object. In Eq. (7), $\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)$ describes an engineering process that requires mathematical representation. The learning process for the input $\mathbf{x}$ primarily consists of two parts: $\mathbf{V}_{1},\ldots,\mathbf{V}_{h}$ and $\mathbf{H}_{1},\ldots,\mathbf{H}_{h}$. $\mathbf{H}_{i}$ is derived based on the parameters $\mathbf{W}_{Qi}$ and $\mathbf{W}_{Ki}$, and can be directly regarded as parameters matrix now, a deep explanation is given in Section 3.5.
@@
Figure 4.a represents the computation process of $\mathbf{H}_{1}\mathbf{V}_{1}=\hat{\mathbf{H}}_{1},\ldots,\mathbf{H}_{h}\mathbf{V}_{h}=\hat{\mathbf{H}}$, and in Figure 4.b, we convert it into the matrix-vector form $\mathbf{W}_{WH}^{\prime}\diamond\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{WH}^{\prime}$ is generated from $\mathbf{H}_{1},\dots,\mathbf{H}_{8}$ and $\mathbf{W}_{V1},\dots,\mathbf{W}_{V8}$. In Figure 4.c, we present a simple example, Figure 4.c.1 depicts $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}]$, while Figure 4.c.2 represents the matrix-vector format of Figure 4.c.1. More ditails could be found in Appendix B.2.
@@
Figure 5 illustrates the parameter transformation scenario after incorporating $\mathbf{W}_{O}$. Figure 5.a demonstrates an example of $\mathring{\mathbf{H}}=\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}$, while Figure 5.b rewrites Figure 5.a as $\mathbf{H}^{\prime}\diamond\mathbf{W}_{O}^{\prime}=\mathring{\mathbf{H}}^{\prime}$. Consequently, based on Figures 4 and 5, the entire MHA can be expressed as $\mathbf{W}_{HVO}^{\prime}\diamond\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{HVO}^{\prime}=\mathbf{W}_{WH}^{T}\diamond\mathbf{W}_{O}^{\prime}$. So the whole MHA can be written as:
$\mathring{\mathbf{H}}$ $=(\mathbf{W}_{HVO}^{\prime})^{T}(\mathbf{x}_{i})^{\prime}$ (8)
@@
Here, $\mathbf{x}_{i}\in\mathbb{R}^{(NM,1)}$ and $\mathring{\mathbf{H}}\in\mathbb{R}^{(NM,1)}$ are the input and output of $i$th layer, respectively, while $\mathbf{W}_{HVO}^{\prime}\in\mathbb{R}^{(NM,NM)}$ denotes matrices generated in accordance with $\mathbf{H}$, $\mathbf{W}_{V}$ and $\mathbf{W}_{O}$. In this manner, we have expressed MHA as a matrix-vector multiplication. This matrix multiplication representation provides a more concise way to express MHA mechanism.
Based on the derivations above and referring to Figure 4, we can gain a deeper understanding of the learning process in MHA. In particular, MHA can be regarded as a specialized form of linear transformation, which involves firstly dividing $\mathbf{x}$ into $\mathbf{x}_{1}...\mathbf{x}_{h}$ based on the number of heads, then transforming $\mathbf{x}_{i}$ into $\mathbf{x}_{i}^{\prime}$, and subsequently utilizing $\mathbf{H}_{i}^{\prime}$, $\mathbf{W}_{Vi}^{\prime}$, and $\mathbf{W}_{Oi}^{\prime}$ to generate the corresponding parameters $\mathbf{W}_{HVOi}^{\prime}$ for $\mathbf{x}_{i}^{\prime}$.
3.5 What makes Transformer to be the winner in LLMs?

### v3

# Universal Approximation Theory: The Basic Theory for Transformer-based Large Language Models

@@ Abstract
Language models have become a focal point in the field of artificial intelligence, especially following the emergence of ChatGPT. Large-scale Transformer networks have quickly become the primary method for advancing natural language processing algorithms. These models, built on the Transformer architecture, are capable of engaging in near-human-like interactions and can even assist in guiding human tasks. Despite their impressive capabilities, there remain some critical theoretical questions in the development of large language models (LLMs): What makes the Transformer architecture so effective in driving intelligent language applications, such as translation and coding? What underlies the in-context learning (ICL) capability of LLMs? How does the LoRA approach enhance the fine-tuning of LLMs? What supports the feasibility of pruning in LLMs? To address these key questions and explore the technical strategies within LLMs, we draw on the Universal Approximation Theory (UAT) to provide a theoretical foundation, revealing the mechanisms behind these advancements.
@@ 1 Introduction
In recent years, the rapid emergence of LLMs in the field of artificial intelligence has undoubtedly become one of the most notable advancements within the domain. The core allure of these models stems from their extraordinary capabilities in language processing. Language, as a unique crystallization of human intelligence, serves not only as the external reflection of thought but also as the bridge for communication, the cornerstone for the dissemination of knowledge, and the continuation of civilization, profoundly shaping the identity of humans as a unique species. Thus, endowing machines with the ability to understand and generate language marks a significant leap towards the realization of true artificial intelligence. The emergence of models such as the ChatGPT (Radford and Narasimhan 2018; Brown et al. 2020; Achiam et al. 2023), the Llama (Touvron et al. 2023), and the PaLM (Chowdhery et al. 2023) vividly demonstrates this point.
@@
A distinctive feature of LLMs is their immense parameter size (Achiam et al. 2023; Touvron et al. 2023; Chowdhery et al. 2023; Chen et al. 2021; Zeng et al. 2022), often amounting to hundreds of billions or even trillions (for instance, GPT-3’s (Brown et al. 2020) 175 billion parameters and PaLM’s (Chowdhery et al. 2023) 540 billion parameters). This vast parameter scale lays the foundation for their exceptional language processing capabilities and enables them to exhibit almost human-like traits, such as ICL  (Brown et al. 2020; Dong et al. 2022), instruction following (Sanh et al. 2021; Ouyang et al. 2022; Wei et al. 2021), and multi-step reasoning (Wei et al. 2022). Notably, these colossal models are predominantly trained by tech giants like Google and Microsoft using large-scale GPU clusters (Zhao et al. 2023), sparking a research fervor on how to efficiently fine-tune them with limited GPU resources. The advent of Lora (Hu et al. 2021) fine-tuning technology has provided an effective pathway for this, allowing for the fine-tuning of large models under resource constraints without the need to adjust all parameters of the original model comprehensively. Moreover, model pruning techniques (Sun et al. 2023; Ma, Fang, and Wang 2023) are crucial for deploying large models in resource-constrained environments, aiming to reduce the model size for operation on smaller devices. Today, LLMs possess a diverse range of functionalities, from translation and text summarization to automatic code generation, demonstrating their versatility.
@@
Despite the rapid advancements of LLMs towards higher intelligence and reliability, the theoretical foundations behind these models remain largely unexplored and shrouded in uncertainty. The scientific community is actively seeking to uncover the intrinsic mechanisms behind their powerful capabilities, including the analysis of ICL mechanisms (Xie et al. 2021; Min et al. 2022). In this context, we establish the UAT (Cybenko 2007; Hornik, Stinchcombe, and White 1989) as the mathematical essence of Transformer-based LLMs and use the perspective of UAT to elucidate key technologies and phenomena within LLMs, aiming to provide explanations from a new angle.
While previous studies have explored the relationship between Transformers and UAT (Yun et al. 2019, 2020; Kratsios et al. 2021; Alberti et al. 2023), these studies are based on a series of complex assumptions and derivations, lacking generality and failing to intuitively address some of the issues present in LLMs. In contrast, this paper aims to provide proof that bridges Transformers and UAT, grounded in the foundational mathematical form of UAT as established by Cybenko; Hornik, Stinchcombe, and White. Our proof is characterized by its simplicity, absence of assumptions, and generality, making it applicable to Convolutional Neural Networks (CNNs) as well (Wang and Li 2024b). Our approach not only provides intuitive answers to some of the challenges within LLMs but also aids in the design of deep learning networks from a theoretical standpoint (Wang and Li 2024a). These capabilities are not addressed by previous theories. Our contributions are as follows:
@@ 2 The Universal Approximation Theory
Up to now, the UAT (Cybenko 2007) is the most widely recognized fundamental theory in deep learning. However, this theorem is only applicable to the simplest form of neural networks, the multilayer perceptron (MLP) (Cybenko 2007; Popescu et al. 2009). Due to the increased complexity of Transformer networks, they cannot be mathematically expressed in the same form as UAT. Thus, the theory has not yet been extended to Transformer networks. Our aim in this paper is to unify Transformer networks under the framework of UAT, thereby standardising their mathematical representation. Before unifying their mathematical forms, we provide a brief overview of the theorem, which was initially proposed by Cybenko (2007). This theorem encompasses numerous conclusions and proof details. Although it has been further developed, its fundamental mathematical form remains unchanged. Therefore, this paper explains the theory based on the UAT form presented by Cybenko (2007). Theorem 2 from Cybenko (2007) states that if $\sigma$ is any continuous sigmoidal function, then finite sums of the following form:
@@
This implies that, when $N$ is sufficiently large, a neural network can approximate any continuous function on a closed interval. Hornik, Stinchcombe, and White (1989) further demonstrates that multilayer feedforward networks also conform to the UAT, capable of approximating arbitrary Borel measurable functions. Observing Equation (1), where the function $G(\mathbf{x})$ yields a scalar output in $\mathbb{R}$, the scenario expands naturally when $G(\mathbf{x})$ maps to $\mathbb{R}^{m}$, requiring the approximation in each dimension. It becomes evident that to accommodate this multidimensional output, a simple extension to Equation (1) suffices: the transformation matrix $\mathbf{W}_{j}$ is revised to the space $\mathbb{R}^{n\times m}$, the bias term $\theta_{j}$ is recast as a vector in $\mathbb{R}^{m}$, and $\alpha_{j}$ is reshaped into a matrix. Nevertheless, in these formulas, the theorem does not straightforwardly apply to Transformer architectures. However, if we succeed in reformulating both the Linear ($\mathbf{x}|\mathbf{W}$) and MHA ($\mathbf{x}|\mathbf{W}$) components into a unified representation, $\mathbf{W}^{\prime}\mathbf{x}^{\prime}$, where $\mathbf{W}$ and $\mathbf{x}$ represent the parameters and inputs for each component respectively, and $\mathbf{W}^{\prime}$ is derived from $\mathbf{W}$ while $\mathbf{x}^{\prime}$ signifies a column vector derived from $\mathbf{x}$ (for details, refer to Section 3.1). It is easy to prove that multi-layer Transformer networks are concrete embodiments of UAT.
@@ 3.2 Diamond Matrix Multiplication
While our goal is to transform Linear and MHA modules into a matrix-vector format, this process is complex. In order to describe this transformation process clearly, we propose a new computational method called the Diamond Multiplication Method, denoted by the symbol $\diamond$. (The purpose of proposing this method is to clearly illustrate the calculation process in turning MHA into Matrix-Vector format. If we use direct matrix multiplication, some computational details cannot be visually distinguished.)
@@
$=\left(\begin{array}[]{c}w_{1}x_{1}+w_{21}x_{2}+\cdots w_{m,1}x_{m}\\ w_{1,2}x_{1}+w_{2,2}x_{2}+\cdots w_{m,2}x_{m}\\ \vdots\\ w_{1,n}x_{1}+w_{2,n}x_{2}+\cdots w_{mn}x_{m}\end{array}\right)$
Let $\mathbf{W}\in\mathbb{R}^{(m,n)}$ and $\mathbf{x}\in\mathbb{R}^{(m,1)}$, the Diamond Matrix Multiplication is defined as $\mathbf{y}=\mathbf{W}\diamond\mathbf{x}$, where $\mathbf{y}\in\mathbb{R}^{(m,1)}$. The computation procedure involves element-wise multiplication and summation of corresponding elements from left to right across the columns of matrix $\mathbf{W}$ with vector $\mathbf{x}$. The sum of the multiplication of the $i$th column of $\mathbf{W}$ with $\mathbf{x}$ is assigned as the $i$th element of $\mathbf{y}$. A detailed calculation process is provided in Eq. (4). The computation process for $\mathbf{y}=\mathbf{x}\diamond\mathbf{W}$ is identical to that of $\mathbf{W}\diamond\mathbf{x}$, implying that $\mathbf{W}\diamond\mathbf{x}=\mathbf{x}\diamond\mathbf{W}$.
@@ 3.3 Matrix-Vector Method for Linear
Here, $\mathbf{x}_{i}\in\mathbb{R}^{(N,M)}$ and $\mathbf{x}_{i+1}\in\mathbb{R}^{(N,M)}$ represent the input and output of layer $i$, respectively, while $\mathbf{W}_{i}\in\mathbb{R}^{(N,N)}$ represents the parameters of layer $i$. $\mathbf{x}_{i}^{\prime}$, $\mathbf{x}_{i+1}^{\prime}$, and $\mathbf{W}_{i}^{\prime}$ are generated based on $\mathbf{x}^{i}$, $\mathbf{x}^{i+1}$, and $\mathbf{W}_{i}$ using the Matrix-Vector Method. For convenience, we use $\mathbf{W}_{i}^{\prime}$ to represent $(\mathbf{W}_{i}^{\prime})^{T}$. Therefore, $\mathbf{x}_{i+1}^{\prime}=\mathbf{W}_{i}^{\prime}\mathbf{x}_{i}^{\prime}$. So the FFN in the Transformer could be written as:
$FFN(\mathbf{x})$ $=\mathbf{W}_{2}\sigma(\mathbf{W}_{1}\mathbf{x}+\mathbf{b}_{1})+\mathbf{b}_{2}$ (6)
$\rightarrow\mathbf{W}_{2}^{\prime}\sigma(\mathbf{W}_{1}^{\prime}\mathbf{x}^{\prime}+\mathbf{b}_{1}^{\prime})+\mathbf{b}_{2}^{\prime}$
@@
Figure 4: The process of transforming $\hat{\mathbf{H}}=\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}}_{8})$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{HV}^{\prime}\diamond\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$.
Figure 5: This diagram illustrates the process of transforming $\text{Concat}(\hat{\mathbf{H}}_{1}...\hat{\mathbf{H}})\mathbf{W}_{O}$ in the MHA into its corresponding matrix-vector form $\mathbf{W}_{HVO}^{\prime}\diamond\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, different colors in the diagram correspond to different variables.
@@ 3.4 Matrix-Vector Method for MHA
$=$ $\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}=\hat{\mathbf{H}}\mathbf{W}_{O}$
@@
Here, $h$ represents the number of attention heads, and the input $\mathbf{x}\in\mathbb{R}^{(N,M)}$ is divided into $\mathbf{x}_{1},...,\mathbf{x}_{h}$ based on $h$. The parameters $\mathbf{W}_{Qi},\mathbf{W}_{Ki}$, and $\mathbf{W}_{Vi}$ correspond to $\mathbf{x}_{i}$. The whole process of MHA can be represented in Figure 3. Figure 3.a represents that the input $\mathbf{x}$ is split into $\mathbf{x}_{1}\cdots\mathbf{x}_{8}$ based on the number of heads. Figure 3.b represents $\mathbf{x}_{i}\mathbf{W}_{Qi}=\mathbf{Q}_{i},\mathbf{x}_{i}\mathbf{W}_{Ki}=\mathbf{K}_{i},\mathbf{x}_{i}\mathbf{W}_{Vi}=\mathbf{V}_{i}$. Figure 3c, d, e represent the process of $\operatorname{softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{T}}{\sqrt{M}}\right)\mathbf{V}_{i}$ and we omit $\sqrt{M}$. Figure 3.f represents $\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}$. Figure 3.g represents the whole process by a matrix multiplication of $(\mathbf{W}^{\prime}_{HVO})^{T}\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, where $\mathbf{W}^{\prime}_{HVO}$ is generated based $\mathbf{H}_{1}\cdots\mathbf{H}_{8}$, $\mathbf{W}_{V1}\cdots\mathbf{W}_{V8}$ and $\mathbf{W}_{O}$. This means that we could represent the whole complex MHA in matrix multiplication. Next, we give the proof of this process.
@@
Our objective is to express the MHA as $(\mathbf{W}^{\prime}_{HVO})^{T}\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$. Prior to transforming the MHA into the matrix-vector form, we need to conduct a comprehensive analysis and clearly define the research object. In Eq. (7), $\hat{\mathbf{H}}=\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)$ describes an engineering process that requires mathematical representation. The learning process for the input $\mathbf{x}$ primarily consists of two parts: $\mathbf{V}_{1},\ldots,\mathbf{V}_{h}$ and $\mathbf{H}_{1},\ldots,\mathbf{H}_{h}$. $\mathbf{H}_{i}$ is derived based on the parameters $\mathbf{x}_{i}$, $\mathbf{W}_{Qi}$ and $\mathbf{W}_{Ki}$.
@@
Figure 4.a represents the computation process of $\mathbf{H}_{1}\mathbf{V}_{1}=\hat{\mathbf{H}}_{1},\ldots,\mathbf{H}_{h}\mathbf{V}_{h}=\hat{\mathbf{H}}_{h}$. In Figure 4.b, we present a simple example of Figure 4.a. Figure 4.b depicts $\mathbf{H}_{i}[\mathbf{x}_{i}\mathbf{W}_{Vi}],i=1,2$. In Figure 4.c we convert Figure 4.b into the matrix-vector form $\mathbf{W}_{HV}^{\prime}\diamond\mathbf{x}^{\prime}=\hat{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{HV}^{\prime}$ is generated from $\mathbf{H}_{1},\mathbf{H}_{2}$ and $\mathbf{W}_{V1},\mathbf{W}_{V2}$. More details can be found in Appendix B.
@@
Figure 5 illustrates the parameter transformation scenario after incorporating $\mathbf{W}_{O}$. Figure 5.a demonstrates an example of $\mathring{\mathbf{H}}=\operatorname{Concat}\left(\hat{\mathbf{H}}_{1},\ldots,\hat{\mathbf{H}}_{h}\right)\mathbf{W}_{O}$, while Figure 5.b rewrites Figure 5.a as $\hat{\mathbf{H}}^{\prime}\diamond\mathbf{W}_{O}^{\prime}=\mathring{\mathbf{H}}^{\prime}$. Consequently, based on Figures 4 and 5, the entire MHA can be expressed as $\mathbf{W}_{HVO}^{\prime}\diamond\mathbf{x}^{\prime}=\mathring{\mathbf{H}}^{\prime}$, where $\mathbf{W}_{HVO}^{\prime}=\mathbf{W}_{HV}^{T}\diamond\mathbf{W}_{O}^{\prime}$ (Based on Figures 5, we can get $\mathbf{W}_{HVO}^{\prime}$, but the resulting expression is too complex and not easy to understand, so we did not compute the final result. However, it is easy to analyze that as long as the variables in $\mathbf{W}_{HV}^{T}$ and $\mathbf{W}_{O}^{\prime}$ are non-zero, the result is essentially a dense matrix). So the whole MHA can be written as:
$\mathring{\mathbf{H}}^{\prime}$ $=(\mathbf{W}_{HVO}^{\prime})^{T}\mathbf{x}^{\prime}$ (9)
@@
Here, $\mathbf{x}\in\mathbb{R}^{(NM,1)}$ and $\mathring{\mathbf{H}}\in\mathbb{R}^{(NM,1)}$ are the input and output, respectively, while $\mathbf{W}_{HVO}^{\prime}\in\mathbb{R}^{(NM,NM)}$ denotes matrices generated in accordance with $\mathbf{H}$, $\mathbf{W}_{V}$ and $\mathbf{W}_{O}$. In this manner, we have expressed MHA as a matrix-vector multiplication. This matrix multiplication representation provides a more concise way to express the MHA mechanism.
For convenience, we use $\mathbf{W}^{\prime}$ to represent $(\mathbf{W}_{HVO}^{\prime})^{T}$. The MHA operation could be simply written as:
$\mathring{\mathbf{H}}^{\prime}=\mathbf{W}^{\prime}\mathbf{x}^{\prime}$ (10)
3.5 The UAT Format of Multi-Layer Transformer